All answers must be entered as a formula !
1. Calculate the operating cash flow and net working capital. Calculating Total Cash Flows Excel ?困- × HOME INSERT PAGE LAYOUT FORMULAS DATA REVIEWVIEW Sign In B 1 u . 田. _a . Alignment Number Cond tional Formatas styles. Cells Editing Formatting Table Clipboard Font Styles Al Jetson Spacecraft Corp. shows the following information on its 2011 income statement: sales $235,000; costs = $141,000; other expenses-$7,900, depreciation expense-S 17,300; interest expense $12,900; taxes $19,565; dividends $12,300. In addition, youre told that the firm issued $6,100 in new equity during 2011 and redeemed $4,500 in outstanding long-term debt. a. What is the 2011 operating cash flow? b. What is the 2011 cash flow to creditors? c. What is the 2011 cash flow to stockholders? d. If net fixed assets increased by $25,000 during the year, what was the addition to NWC? Sales Costs Other expenses Depreciation expense Interest expense Taxes Dividends $ 235,000 41,000 7,900 17,300 12,900 19,565 12,300 10 12 13 14 15 16 2011 New equity Net new long-term debt S 6,100 (4,500) 25,000 17 Change in net fixed assets 18 19 Sheet1 100% Attempt(s)
1. Calculate the operating cash flow and net working capital Calculating Total Cash Flows-Exce FILE HOME INSERTPAGE LAYOUT FORMULAS DATA REVIEW VIEw Sign In 郾 Paste B 1 u . |田·|ク·▲ . Alignment Number Conditional Format Cell Cells Editing FormattingTable Styles Styles AI Income Statement 24 Sales 2s Costs Other expenses Depreciation expense Interest expense Taxes 28 EBIT EBT 2 Net income 4 Dividends 5 Addition to retained earnings 37 38 aOperating cash flow 42 b. Cash flow to creditors 44 C. Cash flow to stockholders d. Cash flow from assets 48 Net capital spending so Change in NWC Sheeti 20%
Answer #1

We need to find OCF first to calculate Net income

Income statement

Sales $235,000

Costs 141,000

Other expenses 7,900

Depreciation 17,300

EBIT $ 68,800

Interest 12,900

Taxable income $ 55,900

Taxes 19,565

Net income $ 36,335

Dividends $12,300

Additions to RE $24,035

a. OCF = EBIT + Depreciation – Taxes = $68,800 + 17,300 – 19,565 = $66,535

b. CFC = Interest – Net new LTD = $12,900 – (–4,500) =$17,400

Note that the net new long-term debt is negative because the company repaid part of its longtermdebt.

c. CFS = Dividends – Net new equity = $12,300 – 6,100 = $6,200

d. We know that CFA = CFC + CFS, so:

CFA = $17,400 + 6,200 = $23,600

CFA is also equal to OCF – Net capital spending – Change in NWC.

We already know OCF.

Netcapital spending is equal to:

Net capital spending = Increase in NFA + Depreciation = $25,000 + 17,300 = $42,300

Now we can use:

CFA = OCF – Net capital spending – Change in NWC

$23,600 = $66,535 – 42,300 – Change in NWC

Change in NWC = $635

This means that the company increased its NWC by $635.